OpenStudy (hatatdatldk):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 3(x + 8)2(x - 8)3

OpenStudy (welshfella):

3(x + 8)^2(x - 8)^3 = 0 so we can say (x + 8)^2 = 0 or (x - 8)^3 = 0

OpenStudy (welshfella):

if (x + 8)^2 = 0 then x + 8 = 0 so x = ?

OpenStudy (hatatdatldk):

Answers: -8, multiplicity 2; 8, multiplicity 3 4, multiplicity 1; 8, multiplicity 1; -8, multiplicity 1 -8, multiplicity 3; 8, multiplicity 2 4, multiplicity 1; -8, multiplicity 3; 8, multiplicity 3

OpenStudy (hatatdatldk):

Ummm 8x?

OpenStudy (welshfella):

x + 8 = 0 you need to isolate x so you subtract 8 from both sides x + 8 - 8 = 0 - 8 so?

OpenStudy (welshfella):

the results will be a number not another term in x

OpenStudy (welshfella):

x = 8x does not make sense

OpenStudy (hatatdatldk):

Yeah..I'm not sure how to do this..sorry

OpenStudy (hatatdatldk):

It would equal -8? @welshfella

OpenStudy (welshfella):

yes -8 is correct x + 8 - 8 = 0 - 8 (8 - 8 = 0) so x = 0-8 = -8 and its multiplicity 2 becaus e of the ^2.

OpenStudy (welshfella):

Now you deal with ( x - 8)^3 in exactly the same way x - 8 = 0 solve for x Multiplicity will be 3

OpenStudy (hatatdatldk):

Okay thank makes sense

OpenStudy (hatatdatldk):

So then I'd say it's A?

OpenStudy (welshfella):

you isolate x this time by adding 6 to both sides

OpenStudy (welshfella):

* adding 8

OpenStudy (welshfella):

yea

OpenStudy (welshfella):

x = 8 for the (x - 8)^3 = 0

OpenStudy (hatatdatldk):

Yes, thank you